### v1

# Stability of the Kim–Milman flow map

Abstract
In this short note, we characterize stability of the Kim–Milman flow map—also known as the probability flow ODE—with respect to variations in the target measure. Rather than the Wasserstein distance, we show that stability holds with respect to the relative Fisher information.
1 Introduction
In general, there are infinitely many maps which transport a fixed source distribution $\rho$ to a target distribution $\mu$, both in $\mathcal{P}(\mathbb{R}^{d})$. Let $T_{\bullet}$ denote a method of generating such transport maps; thus, $T_{\bullet}^{\rho\to\mu}$ is a transport map from $\rho$ to $\mu$ for every $\mu\in\mathcal{P}(\mathbb{R}^{d})$, meaning that for $X\sim\rho$, $T_{\bullet}^{\rho\to\mu}(X)\sim\mu$. A property of fundamental interest for such a method is its stability with respect to variations in the target measure: how much do the transport maps vary if the target measures vary? That is, for another target measure $\nu$ and under a broad class of assumptions, we want to understand inequalities of the form
$\|T^{\rho\to\mu}_{\bullet}-T_{\bullet}^{\rho\to\nu}\|^{2}_{L^{2}(\rho)}\lesssim\mathrm{D}(\mu,\nu)\,$, (1.1)
where $\rm{D}(\mu,\nu)$ is some dissimilarity metric between the two target measures.
To the best of our knowledge, the study of inequalities of the form (1.1) has been limited to optimal transport maps [11, 14, 17, 2], denoted $T_{\rm OT}^{\rho\to\mu}$, or entropic transport maps [4, 9], denoted $T_{\rm EOT}^{\rho\to\mu}$. In these instances, the natural dissimilarity metric becomes the (squared) $2$-Wasserstein distance between $\mu$ and $\nu$, and the underlying constant depends on properties of the source $\rho$, and either the class of target measures $\mu$ or a priori assumptions on the (entropic) optimal transport map. Existing bounds are of the form
$\|T^{\rho\to\mu}_{\rm OT}-T_{\rm OT}^{\rho\to\nu}\|_{L^{2}(\rho)}\leq CW_{2}^{{\beta}}(\mu,\nu)\,,\quad\text{ or }\quad\|T^{\rho\to\mu}_{\rm EOT}-T_{\rm EOT}^{\rho\to\nu}\|_{L^{2}(\rho)}\leq C_{\varepsilon}W_{2}(\mu,\nu)\,$,
where $\beta\in(0,1]$ and, for the entropic maps, $\varepsilon>0$ is the regularization parameter and $C_{\varepsilon}\nearrow+\infty$ as $\varepsilon\searrow 0$; see Section 2.2 for more information.
In this note, we study stability properties of a different transport map called the Kim–Milman (reverse) heat flow map [13], which has recently gained popularity in the machine learning literature under the moniker “probability flow ODE” [12] due to its efficacy in generative modeling tasks. Unlike optimal or entropic transport maps, we stress that this transport map is defined dynamically. As a brief description, let $\gamma$ denote the $d$-dimensional standard Gaussian distribution, and let $\mu$ be another $d$-dimensional probability distribution; we want to study the stability of the following (reverse) ODE system
$\dot{X}_{t}$ $=X_{t}+\nabla\log\mu Q_{T-t}(X_{t})=\nabla\log Q_{T-t}\Bigl[\frac{\mu}{\gamma}\Bigr](X_{t})\,$,
where $0\ll T<\infty$, $Q_{s}$ is the Ornstein–Uhlenbeck semigroup at time $s>0$, and we initialize $X_{0}\sim\mu Q_{T}\approx\gamma$. Let $T_{\rm KM}^{\mu}$ be the flow map of the ODE system with $T=\infty$ (rigorously, the limit of the flow maps up to time $T$, taking $T\nearrow\infty$), which transports $(T_{\rm KM}^{\mu})_{\sharp}\gamma=\mu$. By imposing regularity assumptions on $\mu$, we will prove that
$\|T^{\mu}_{\rm KM}-T_{\rm KM}^{\nu}\|_{L^{2}(\gamma)}\lesssim\sqrt{\rm{FI}(\nu\,\|\,\mu)}\,$, (1.2)
where $\rm{FI}(\nu\,\|\,\mu)\coloneqq\|\nabla\log(\nu/\mu)\|^{2}_{L^{2}(\nu)}$, where the underlying constant is explicit. We then show how our analysis can be used to prove bounds of the form
$\|T^{\mu}_{\rm KM}-T_{\rm KM}^{\nu}\|_{L^{\infty}(\gamma)}\lesssim\sqrt{\operatorname{FI}_{\infty}(\nu\,\|\,\mu)}\,$, (1.3)
where $\operatorname{FI}_{\infty}(\nu\,\|\,\mu)\coloneqq{\rm{ess\,sup}}_{\nu}\|\nabla\log(\nu/\mu)\|^{2}$. We consider as applications the case where $\mu$ is a perturbation of a strongly log-concave measure (leveraging recent results by [3, 26]) and when $\mu$ has an asymptotically positive convexity profile (as introduced by [7]). We stress that, at present, these cases are not covered by stability results for (entropic) optimal transport maps.
Notation
We write $\mathcal{P}(\mathbb{R}^{d})$ for the space of probability measures over $\mathbb{R}^{d}$. Given a probability measure $\mu$ which admits a Lebesgue density, we abuse notation and write $\mu$ for the density as well as for the measure. For a symmetric matrix $\Sigma$, we write $\|x\|_{\Sigma}\coloneqq\sqrt{\langle x,\Sigma\,x\rangle}$.
2 Background
2.1 Probability flow ODE
We now review the derivation of the probability flow ODE, also known as the reverse heat flow due to Kim and Milman [13].
For ${X}^{\rightarrow}_{0}\sim\mu\in\mathcal{P}(\mathbb{R}^{d})$, recall the forward Ornstein–Uhlenbeck process
${\rm d}{X}^{\rightarrow}_{t}=-{X}^{\rightarrow}_{t}{\rm d}t+\sqrt{2}\,{\rm d}B_{t}\,$,
where $(B_{t})_{t\geq 0}$ is standard Brownian motion. Note that as $t\to\infty$, $q_{t}\coloneqq{\rm Law}(X_{t}^{\rightarrow})\to\gamma=\mathcal{N}(0,I)$. Now, run the stochastic differential equation (SDE) for time $0\ll T<+\infty$. Then, the reverse SDE system is given by
${\rm d}{X}^{\leftarrow}_{t}=\bigl({X}^{\leftarrow}_{t}+2\,\nabla\log q_{T-t}({X}^{\leftarrow}_{t})\bigr)\,{\rm d}t+\sqrt{2}\,{\rm d}B_{t}\,,\qquad X_{0}^{\leftarrow}\sim q_{T}\,$, (2.1)
where ${\rm Law}(X^{\leftarrow}_{s})=q_{T-s}$. (Precise conditions for the well-posedness of the time reversal can be found in [5].) Note that the Brownian motion is also reversed. The corresponding Fokker–Planck equation for the reverse SDE is then
$\partial_{t}q_{T-t}+\nabla\cdot(q_{T-t}\,({\rm{id}}+2\,\nabla\log q_{T-t}))=\Delta q_{T-t}=\nabla\cdot(q_{T-t}\nabla\log q_{T-t})\,.$
We can incorporate the diffusion term above into the drift, resulting in the continuity equation
$\partial_{t}q_{T-t}+\nabla\cdot(q_{T-t}\,({\rm{id}}+\nabla\log q_{T-t}))=0\,$,
which describes the evolution of marginal law of the ODE system
$\dot{X}^{\leftarrow}_{t}={X}^{\leftarrow}_{t}+\nabla\log q_{T-t}({X}^{\leftarrow}_{t})\,,\qquad X_{0}^{\leftarrow}\sim q_{T}\,.$ (2.2)
Note that while (2.1) and (2.2) differ as stochastic processes, by construction they have the same marginal laws $(q_{T-t})_{t\in[0,T]}$.
Let $T_{\rm KM}^{\mu,T}$ be the flow map corresponding the ODE system (2.2); thus, $(T_{\rm KM}^{\mu,T})_{\sharp}q_{T}=\mu$. Finally, we let $T_{\rm KM}^{\mu}\coloneqq\lim_{T\to\infty}T_{\rm KM}^{\mu,T}$ denote the Kim–Milman map.
2.2 Related work
Stability of optimal transport maps.
Stability of optimal transport maps was first studied by Gigli in [11]. His main result states that if one of the transport maps, say $T^{\rho\to\mu}_{\rm OT}$, is $L$-Lipschitz and the support of $\mu$ is compact (say in a ball with radius $R$), then
$\|T^{\rho\to\mu}_{\rm OT}-T_{\rm OT}^{\rho\to\nu}\|_{L^{2}(\rho)}\lesssim(LR)^{1/2}\,W_{2}^{1/2}(\mu,\nu)\,.$
Most recently, this argument has been pushed by Letrouit and Mérigot in [17] to show that, so long as $\mu,\nu$ have compact support and $\rho$ has density bounded above and below on a compact convex subset of $\mathbb{R}^{d}$, then
$\|T^{\rho\to\mu}_{\rm OT}-T_{\rm OT}^{\rho\to\nu}\|_{L^{2}(\rho)}\lesssim W_{2}^{1/6}(\mu,\nu)\,.$
We note that this same result is true over Riemannian manifolds [14]. On the other hand, it is not possible to establish Hölder stability in general with an exponent better than $1/2$, due to the counterexample in [11].
Going a step further, Manole, Balakrishnan, Niles-Weed, and Wasserman [19] show that if one of the optimal transport maps is bi-Lipschitz, i.e., if $0\prec\ell I\preceq DT^{\rho\to\mu}_{\rm OT}\preceq LI$, then we have the stronger bound
$\|T^{\rho\to\mu}_{\rm OT}-T_{\rm OT}^{\rho\to\nu}\|_{L^{2}(\rho)}\leq\Bigl(\frac{L}{\ell}\Bigr)^{1/2}\,W_{2}(\mu,\nu)\,.$
For more results on this topic, see the recent monograph by Letrouit [16].
Stability of entropic transport maps.
Entropic transport maps were recently introduced in [21] to estimate optimal transport maps from samples; see also [24, 23]. The stability of entropic transport maps is more recent, first investigated by Carlier, Chizat, and Laborde [4]. Specializing their results, they prove that if all measures $\rho,\mu,\nu$ have compact support, then
$\|T^{\rho\to\mu}_{\rm EOT}-T_{\rm EOT}^{\rho\to\nu}\|_{L^{2}(\rho)}\lesssim\exp(c/\varepsilon)\,W_{2}(\mu,\nu)\,$,
where $c>0$ is a constant depending on the diameter of the supports. Crucially, as $\varepsilon\searrow 0$, the results for optimal transport maps are not recovered in any regime. Recently, a tight stability bound for entropic transport maps was proven by Divol, Niles-Weed, and the second author [9], showing that
$\|T^{\rho\to\mu}_{\rm EOT}-T_{\rm EOT}^{\rho\to\nu}\|_{L^{2}(\rho)}\lesssim\frac{R^{2}}{\varepsilon}\,W_{2}(\mu,\nu)\,$,
where the sole requirement is that all three measures have compact support (say in a ball of radius $R>0$). Though, the results of [9] are more general. For example, if one of the entropic transport maps is bi-Lipschitz (which can be ensured in certain situations [6]), then they are able to recover the results of [19] by taking $\varepsilon\searrow 0$.
3 Main results
We are interested in understanding the dynamics of the system (2.2). To this end, let $(X_{t})_{t\in[0,T]}$ (resp. $(Y_{t})_{t\in[0,T]}$) be the reverse dynamics from $\gamma$ to $\mu$ (resp. $\gamma$ to $\nu$) for $0\ll T<+\infty$ with $q^{\mu}_{t}\coloneqq{\rm{Law}}(X_{t})$ (resp. $q^{\nu}_{t}\coloneqq{\rm{Law}}(Y_{t})$). Thus, $T_{\rm KM}^{\mu}(x)$ is the terminal point $X_{T}$ of the ODE (2.2) with $T\to\infty$ and initialized at $x$, and similarly for $T_{\rm KM}^{\nu}(x)$. For a rigorous justification of this procedure, see, e.g., [13, 20]. Our calculations will require that for all $s\geq 0$
$\nabla^{2}\log Q_{s}\Bigl[\frac{\mu}{\gamma}\Bigr](\cdot)=I+\nabla^{2}\log q_{s}^{\mu}(\cdot)\preceq{\theta}_{s}I\,.$ ($\Theta$)
Our main results will be of the form
$\|T_{\rm KM}^{\mu}-T_{\rm KM}^{\nu}\|_{L^{2}(\gamma)}\lesssim\sqrt{{\rm FI}(\nu\,\|\,\mu)}\,$,
where we place assumptions on $\mu$, and the omitted constant will depend explicitly on these assumptions. We can view this result as a strengthening of the classical transport-information inequality. For instance, an application of the HWI inequality (see [10]) for $\mu$ which is $\alpha$-strongly log-concave yields
$W_{2}^{2}(\mu,\nu)\leq\alpha^{-2}\,{\rm FI}(\nu\,\|\,\mu)\,$, (3.1)
where $W_{2}^{2}(\mu,\nu)$ denotes the squared $2$-Wasserstein distance between $\mu$ and $\nu$. However, it follows from a trivial coupling argument and Corollary 3.4 that
$W_{2}^{2}(\mu,\nu)\leq\|T_{\rm{KM}}^{\mu}-T_{\rm{KM}}^{\nu}\|^{2}_{L^{2}(\gamma)}\leq\alpha^{-2}\,{\rm FI}(\nu\,\|\,\mu)\,.$
Thus, we have strengthened (3.1) by giving an explicit coupling, and by replacing the $W_{2}$ metric on the left-hand side with a larger quantity (in fact, the “linearized” Wasserstein metric at $\gamma$ between $T_{\rm{KM}}^{\mu}$ and $T_{\rm{KM}}^{\nu}$).
3.1 Main computation
In this section, we prove the following proposition.
Proposition 3.1.
Let $\mu$ be such that ($\Theta$) is satisfied for some constants $(\theta_{s})_{s\geq 0}$. Then, for any $\nu$ such that $\rm{FI}(\nu\,\|\,\mu)<\infty$ and $T\geq 0$,
$(\mathbb{E}\|X_{T}-Y_{T}\|^{2})^{1/2}$ $\leq\sqrt{{\rm FI}(\nu\,\|\,\mu)}\,\Lambda_{T}\coloneqq\sqrt{{\rm FI}(\nu\,\|\,\mu)}\int_{0}^{T}\exp\Bigl(\int_{0}^{T-s}(3\theta_{u}-1)\,\mathrm{d}u\Bigr)\,\mathrm{d}s\,.$
Our main theorem is then the following, which follows as an immediate corollary of Proposition 3.1 by taking the appropriate limits.
Theorem 3.2.
Suppose the conditions of Proposition 3.1 are satisfied, and write $\Lambda_{\infty}\coloneqq\lim_{T\to\infty}\Lambda_{T}$. Then,
$\|T_{\rm KM}^{\mu}-T_{\rm KM}^{\nu}\|_{L^{2}(\gamma)}\leq\Lambda_{\infty}\sqrt{{\rm FI}(\nu\,\|\,\mu)}\,.$
To prove our main proposition, we require the following lemma due to Wibisono [26].
Lemma 3.3 ([26, Theorem 4(i)]).
For any $t>0$,
$\frac{{\rm d}}{{\rm d}t}\,{\rm FI}(q_{t}^{\nu}\,\|\,q_{t}^{\mu})\leq-2\,\mathbb{E}_{q_{t}^{\nu}}\Bigl\|\nabla\log\frac{q^{\nu}_{t}}{q_{t}^{\mu}}\Bigr\|^{2}_{(-2\nabla^{2}\log q_{t}^{\mu}{-}I)}\,.$
Proof. [Proof of Proposition 3.1] To start, we compute
$\partial_{t}\mathbb{E}\|X_{t}-Y_{t}\|^{2}$ $=2\,\mathbb{E}\langle X_{t}-Y_{t},\dot{X}_{t}-\dot{Y}_{t}\rangle$
$=2\,\mathbb{E}\Bigl\langle X_{t}-Y_{t},\nabla\log Q_{T-t}\Bigl[\frac{\mu}{\gamma}\Bigr](X_{t})-\nabla\log Q_{T-t}\Bigl[\frac{\nu}{\gamma}\Bigr](Y_{t})\Bigr\rangle$
$\leq 2\theta_{T-t}\,\mathbb{E}\|X_{t}-Y_{t}\|^{2}$
$\qquad{}+2\,\mathbb{E}\Bigl\langle X_{t}-Y_{t},\nabla\log Q_{T-t}\Bigl[\frac{\mu}{\gamma}\Bigr](Y_{t})-\nabla\log Q_{T-t}\Bigl[\frac{\nu}{\gamma}\Bigr](Y_{t})\Bigr\rangle\,$,
where we used ($\Theta$) in the last inequality. Using Cauchy–Schwarz, we obtain
$\partial_{t}\mathbb{E}\|X_{t}-Y_{t}\|^{2}$ $\leq 2{\theta_{T-t}}\,\mathbb{E}\|X_{t}-Y_{t}\|^{2}+2\,(\mathbb{E}\|X_{t}-Y_{t}\|^{2})^{1/2}\,{\rm{FI}}(q^{\nu}_{T-t}\,\|\,q^{\mu}_{T-t})^{1/2}\,$, (3.2)
where the relative Fisher information makes an appearance. Now, as $\nabla^{2}\log q_{t}^{\mu}=\nabla^{2}\log Q_{t}\bigl[\frac{\mu}{\gamma}\bigr]-I\preceq{(\theta_{t}-1)}\,I$, Lemma 3.3 simplifies to
$\frac{{\rm d}}{{\rm d}t}{\rm FI}(q_{t}^{\nu}\,\|\,q_{t}^{\mu})\leq-2\,(-2\,(\theta_{t}-{1})-1)\,{\rm FI}(q_{t}^{\nu}\,\|\,q_{t}^{\mu})=-2\,(-2\theta_{t}+{1})\,{\rm FI}(q_{t}^{\nu}\,\|\,q_{t}^{\mu})\,$,
via ($\Theta$) and thus, by Grönwall’s inequality
${\rm FI}(q_{t}^{\nu}\,\|\,q_{t}^{\mu})$ $\leq\exp\Bigl(-2\int_{0}^{t}(-2\theta_{u}+{1})\,\mathrm{d}u\Bigr)\,{\rm FI}(q_{0}^{\nu}\,\|\,q_{0}^{\mu})$
$=\exp\Bigl(-2\int_{0}^{t}(-2\theta_{u}+{1})\,\mathrm{d}u\Bigr)\,{\rm FI}(\nu\,\|\,\mu)\,.$
Since the above holds for any time $t>0$, we choose TtT-t, and thus
$\sqrt{{\rm FI}(q_{T-t}^{\nu}\,\|\,q_{T-t}^{\mu})}\leq e^{-(T-t)}\exp\Bigl(\int_{0}^{T-t}2\theta_{u}\,\mathrm{d}u\Bigr)\,{\rm FI}(\nu\,\|\,\mu)^{1/2}\eqqcolon{c_{T-t}}\sqrt{{\rm FI}(\nu\,\|\,\mu)}\,.$
Using the fact that
$\partial_{t}(\mathbb{E}\|X_{t}-Y_{t}\|^{2})^{1/2}=\frac{1}{2}\,\frac{\partial_{t}\mathbb{E}\|X_{t}-Y_{t}\|^{2}}{(\mathbb{E}\|X_{t}-Y_{t}\|^{2})^{1/2}}\,$,
we obtain
$\partial_{t}(\mathbb{E}\|X_{t}-Y_{t}\|^{2})^{1/2}\leq{\theta_{T-t}}\,(\mathbb{E}\|X_{t}-Y_{t}\|^{2})^{1/2}+c_{T-t}\sqrt{{\rm FI}(\nu\,\|\,\mu)}\,.$
Applying Grönwall’s inequality again gives
$(\mathbb{E}\|X_{t}-Y_{t}\|^{2})^{1/2}\leq\sqrt{{\rm FI}(\nu\,\|\,\mu)}\int_{0}^{t}c_{T-s}\exp\Bigl(\int_{s}^{t}{\theta_{T-r}}\,\mathrm{d}r\Bigr)\,\mathrm{d}s\,.$
Taking $t=T$ and writing out $c_{T-s}$, our full inequality is
$(\mathbb{E}\|X_{T}-Y_{T}\|^{2})^{1/2}$ $\leq\sqrt{{\rm FI}(\nu\,\|\,\mu)}\int_{0}^{T}e^{-(T-s)}\exp\Bigl(\int_{0}^{T-s}2\theta_{u}\,\mathrm{d}u\Bigr)\exp\Bigl(\int_{s}^{T}{\theta_{T-r}}\,\mathrm{d}r\Bigr)\,\mathrm{d}s$
$=\sqrt{{\rm FI}(\nu\,\|\,\mu)}\int_{0}^{T}e^{-(T-s)}\exp\Bigl(3\int_{0}^{T-s}\theta_{u}\,\mathrm{d}u\Bigr)\,\mathrm{d}s\,$,
where the final step follows from change of variables. ∎
3.2 Corollaries
We now instantiate Proposition 3.1 with some cases of interest.
3.2.1 Strong log-concavity
As a warm-up, suppose that $\mu$ is $\alpha$-strongly log-concave, i.e., there exists $\alpha>0$ such that $0\prec\alpha I\preceq-\nabla^{2}\log\mu$. In this case, it is well-known that (see, e.g., [20])
${\theta}_{u}=\frac{1-\alpha}{\alpha\,(\exp(2u)-1)+1}\,.$
Incorporating this bound into Proposition 3.1 and carrying out the integration yields the following corollary. As this is a special case of the next section, we omit the computation.
Corollary 3.4.
Suppose that $\mu$ is $\alpha$-strongly convex. Then
$\|T_{\rm{KM}}^{\mu}-T_{\rm{KM}}^{\nu}\|_{L^{2}(\gamma)}\leq\alpha^{-1}\,\sqrt{{\rm FI}(\nu\,\|\,\mu)}\,.$
3.2.2 Strong log-concavity with log-Lipschitz perturbations
We now suppose our main target measure is of the form $\mu\propto\exp(-V+H)$ where $V$ is $\alpha$-strongly convex and $H$ is a (smooth) $L$-Lipschitz perturbation. In this setting, a recent result of [3] showed that
$\theta_{u}=\frac{1-\alpha}{\alpha\,(e^{2u}-1)+1}+\frac{e^{2u}L^{2}}{(\alpha\,(e^{2u}-1)+1)^{2}}+\frac{2Le^{2u}}{(\alpha\,(e^{2u}-1)+1)^{3/2}\sqrt{e^{2u}-1}}\,.$ (3.3)
Carrying out the algebra, the complete stability bound in this setting is given below. Note that the special case $L=0$ recovers Corollary 3.4.
Corollary 3.5 (Perturbation of strongly log-concave).
Suppose $\mu\propto\exp(-V-H)$ where $V$ is $\alpha$-strongly convex and $H$ is an $L$-Lipschitz function. Then
$\|T_{\rm{KM}}^{\mu}-T_{\rm{KM}}^{\nu}\|_{L^{2}(\gamma)}\leq\frac{1}{\alpha}\exp\Bigl(\frac{3L^{2}}{2\alpha}+\frac{6L}{\sqrt{\alpha}}\Bigr)\,\sqrt{{\rm FI}(\nu\,\|\,\mu)}\,.$
Proof.  Letting $b=\exp(2(T-s))-1$, carrying out the integration yields
$\int_{0}^{T-s}\theta_{u}\,\mathrm{d}u=-\frac{1}{2}\log\Bigl(\frac{1+\alpha b}{1+b}\Bigr)+\frac{bL^{2}}{2\,(1+\alpha b)}+\frac{2bL}{1+\alpha b}\sqrt{\alpha+b^{-1}}\,.$
Another change of variables with $r=(b+1)^{-1/2}=\exp(-(T-s))$ yields, for the full integral,
$\int_{0}^{T}\exp(-(T-s))\exp\Bigl(3\int_{0}^{T-s}\theta_{u}\,\mathrm{d}u\Bigr)\,\mathrm{d}s$
$\qquad\qquad=\int_{\exp(-T)}^{1}(r^{2}+\alpha\,(1-r^{2}))^{-3/2}\exp\Bigl(\frac{3L^{2}}{2\,(f(r)+\alpha)}\Bigr)\exp\Bigl(\frac{6L}{\sqrt{f(r)+\alpha}}\Bigr)\,\mathrm{d}r\,$,
where $f(r)=r^{2}/(1-r^{2})$. As $f$ is increasing on the interval $(0,1)$ and $f(0)=0$, we can replace $f(r)+\alpha\geq\alpha$, and obtain
$\int_{0}^{T}\exp(-(T-s))\exp\Bigl(3\int_{0}^{T-s}\theta_{u}\,\mathrm{d}u\Bigr)\,\mathrm{d}s$
$\qquad\qquad\leq\exp\Bigl(\frac{3L^{2}}{2\alpha}+\frac{6L}{\sqrt{\alpha}}\Bigr)I(T,\alpha)$
$\qquad\qquad\coloneqq\exp\Bigl(\frac{3L^{2}}{2\alpha}+\frac{6L}{\sqrt{\alpha}}\Bigr)\int_{\exp(-T)}^{1}(r^{2}+\alpha\,(1-r^{2}))^{-3/2}\,\mathrm{d}r\,.$
Performing the integration in closed form, it is easy to see that $\lim_{T\to\infty}I(T,\alpha)=\alpha^{-1}$. This concludes the proof, as
$\int_{0}^{T}\exp(-(T-s))\exp\Bigl(3\int_{0}^{T-s}\theta_{u}\,\mathrm{d}u\Bigr)\,\mathrm{d}s$ $\leq\frac{1}{{\alpha}}\exp\Bigl(\frac{3L^{2}}{2\alpha}+\frac{6L}{\sqrt{\alpha}}\Bigr)\,.$
Example: Gaussian mixtures as tilts. As an example, we take the case of Gaussian mixtures. Suppose $\mu=\sum_{k=1}^{K}w_{k}\varphi(\cdot;m_{k},\Sigma)$ where $\varphi(\cdot;m_{k},\Sigma)$ is the Gaussian density with mean $m_{k}\in\mathbb{R}^{d}$ and covariance $\Sigma\succ 0$, and $w_{k}\geq 0$ are weights (such that $\sum_{k=1}^{K}w_{k}=1$). In this case, it is possible to write down the log-density in the form of our assumptions, with
$V(x)=\tfrac{1}{2}\|x\|_{\Sigma^{-1}}^{2}\,,\quad H(x)=\log\sum_{k=1}^{K}w_{k}\exp\bigl(m_{k}^{\top}\Sigma^{-1}x-\tfrac{1}{2}m_{k}^{\top}\Sigma^{-1}m_{k}\bigr)\,$,
and, moreover, it is easy to verify that
$\nabla H(x)=\Sigma^{-1}\sum_{k=1}^{K}{w}_{k}(x)\,m_{k}\coloneqq\Sigma^{-1}\frac{\sum_{k=1}^{K}w_{k}m_{k}\exp\bigl(m_{k}^{\top}\Sigma^{-1}x-\tfrac{1}{2}m_{k}^{\top}\Sigma^{-1}m_{k}\bigr)}{\sum_{k=1}^{K}w_{k}\exp\bigl(m_{k}^{\top}\Sigma^{-1}x-\tfrac{1}{2}m_{k}^{\top}\Sigma^{-1}m_{k}\bigr)}\,.$
If we further assume $\alpha I\preceq\Sigma^{-1}\preceq\beta I$, then via Jensen’s inequality,
$\|\nabla H(x)\|\leq\|\Sigma^{-1}\|_{\rm op}\max_{k}\|m_{k}\|\leq\beta\max_{k}\|m_{k}\|\,.$
Thus, for any $\nu$, our stability bound reads
$\|T_{\rm{KM}}^{\mu}-T_{\rm{KM}}^{\nu}\|_{L^{2}(\gamma)}$ $\leq\frac{1}{\alpha}\exp\Bigl(C\,\frac{\beta^{2}\max_{k}\|m_{k}\|^{2}}{\alpha}\Bigr)\,\sqrt{{\rm FI}(\nu\,\|\,\mu)}\,$,
for some universal constant $C>0$.
3.2.3 Distributions with asymptotically positive convex profiles
As a final example, we turn to a family of distributions introduced in [8, 7]. To define said family, we require the following definitions. For a function $f:\mathbb{R}^{d}\to\mathbb{R}$, we define the integrated convexity profile of $f$, denoted $\kappa_{f}:\mathbb{R}_{+}\to\mathbb{R}\cup\{-\infty\}$, to be
$\kappa_{f}(r)\coloneqq\inf\Bigl\{\frac{\langle\nabla U(x)-\nabla U(y),x-y\rangle}{\|x-y\|^{2}}\ :\ \|x-y\|=r\Bigr\}\,.$
We introduce this definition with the following example.
Example: Strongly convex potential outside a ball. Taking $\mu\propto\exp(-V)$, suppose that there exists $\alpha_{V}>0$ and $R_{V},L_{V}\geq 0$ such that
$\kappa_{V}(r)\geq\begin{cases}\alpha_{V}&{\text{ for }}r>R_{V}\,,\\ \alpha_{V}-L_{V}&{\text{ for }}r\leq R_{V}\,.\end{cases}$ (3.4)
If $R_{V}=0$, then $\kappa_{V}(r)\geq\alpha_{V}>0$ is equivalent to strong convexity over all of $\mathbb{R}^{d}$. Otherwise for $R_{V}>0$, the potential has integrated convexity profile which might be negative inside $B(0,R_{V})$, and while remaining strongly convex outside the ball.
The following proposition gives an alternative characterization of (3.4).
Proposition 3.6 ([7, Proposition 5.1]).
Suppose $V$ satisfies (3.4) with constants $\alpha_{V}>0$ and $L_{V},R_{V}\geq 0$. Then it satisfies, for all $r>0$,
$\kappa_{V}(r)\geq\alpha_{V}-r^{-1}\hat{g}_{\hat{L}}(r)\,$,
where $\hat{g}_{L}(r)\coloneqq 2\sqrt{L}\tanh(r\sqrt{L})$, and $\hat{L}$ is given by
$\hat{L}\coloneqq\begin{cases}\inf\{L\geq 0:R_{V}^{-1}\hat{g}_{L}(R_{V})\geq L_{V}\}\,,&R_{V}>0\,,\\ 0\,,&R_{V}=0\,.\end{cases}$
Following Proposition 3.6, it is worth considering some asymptotic scenarios for determining $\hat{L}$. For instance, if $R_{V}^{2}L_{V}\ll 1$, then one can verify that $\hat{L}\approx L_{V}/2$. On the other hand, if $R_{V}^{2}L_{V}\gg 1$, then $\hat{L}\approx L_{V}^{2}R_{V}^{2}/4$.
To generalize this characterization, we can consider functions $g\in\mathcal{G}\subset C^{2}((0,\infty),\mathbb{R}_{+})$ if they satisfy the following properties:
  1. 1.
    $r\mapsto r^{1/2}g(r^{1/2})$ is non-decreasing and concave,
  2. 2.
    $\lim_{r\downarrow 0}rg(r)=0$,
  3. 3.
    $g$ itself is bounded such that $g^{\prime}\geq 0$ and $2g^{\prime\prime}+gg^{\prime}\leq 0$,
  4. 4.
    the right-derivative of $\hat{g}$ at the origin exists, denoted $\hat{g}^{\prime}(0)$.
This function class leads to our final corollary.
Corollary 3.7 (Asymptotically positive convex profile).
Let ${\alpha}>-1$ and suppose $\hat{g}\in\mathcal{G}$. Suppose $\mu\propto\gamma\exp(-h)$ where the integrated convexity profile of $h$ satisfies $\kappa_{h}(r)\geq\alpha-r^{-1}\hat{g}(r)$. Then,
$\|T_{\rm{KM}}^{\mu}-T_{\rm{KM}}^{\nu}\|_{L^{2}(\gamma)}\leq\frac{1}{1+\alpha}\exp\Bigl(\frac{3\hat{g}^{\prime}(0)}{2\,(1+\alpha)}\Bigr)\,\sqrt{{\rm FI}(\nu\,\|\,\mu)}\,.$
Proof.  By [8, Lemma 5.9], in this case it holds that
$\theta_{u}=-\frac{\exp(-2u)}{1+(1-\exp(-2u))\,\alpha}\Bigl(\alpha-\frac{\hat{g}^{\prime}(0)}{1+(1-\exp(-2u))\,\alpha}\Bigr)\,.$
Mimicking the computations from before, we then compute
$\int_{0}^{T}\exp(-(T-s))\exp\Bigl(3\int_{0}^{T-s}\theta_{u}\,\mathrm{d}u\Bigr)\,\mathrm{d}s$
$\qquad\qquad=\int_{\exp(-T)}^{1}(1+\alpha-\alpha r^{2})^{-3/2}\exp\Bigl(-\frac{3\hat{g}^{\prime}(0)}{2\alpha}\bigl((1+\alpha-\alpha r^{2})^{-1}-1)\bigr)\Bigr)\,\mathrm{d}r$
$\qquad\qquad\leq\exp\Bigl(-\frac{3\hat{g}^{\prime}(0)}{2\alpha}\bigl((1+\alpha)^{-1}-1)\bigr)\Bigr)\int_{\exp(-T)}^{1}(1+\alpha-\alpha r^{2})^{-3/2}\,\mathrm{d}r\,$,
where we obtain the result by taking the $T\to\infty$ limit. ∎
Returning to our example of log-densities which are strongly convex outside a ball, we can instantiate Corollary 3.7 with the heuristic bounds on $\hat{L}$ to obtain the bounds
$\|T_{\rm{KM}}^{\mu}-T_{\rm{KM}}^{\nu}\|_{L^{2}(\gamma)}\leq\alpha^{-1}\exp\bigl(O\bigl((L_{V}/\alpha)\,(1\vee L_{V}R_{V}^{2})\bigr)\bigr)\,\sqrt{{\rm FI}(\nu\,\|\,\mu)}\,.$
4 Extension to stronger metrics
In this section, we show how our proof above can be modified to provide stability bounds under stronger metrics. Writing $\operatorname{FI}_{\infty}(\nu\,\|\,\mu)\coloneqq{\|\nabla\log(\nu/\mu)\|_{L^{\infty}(\nu)}^{2}}$, our goal now is to establish bounds of the form
$\|T^{\mu}_{\rm KM}-T^{\nu}_{\rm KM}\|_{L^{\infty}(\gamma)}\lesssim\sqrt{\operatorname{FI}_{\infty}(\nu\,\|\,\mu)}\,$, (4.1)
where again the hidden constants will be made explicit.
To start, we follow the start of the proof of Proposition 3.1 assuming ($\Theta$). It is easy to see that $\gamma$-almost surely, it holds that
$\partial_{t}\|X_{t}-Y_{t}\|^{2}\leq 2\theta_{T-t}\,\|X_{t}-Y_{t}\|^{2}+2\,\|X_{t}-Y_{t}\|\operatorname{FI}_{\infty}(q_{T-t}^{\nu}\,\|\,q_{T-t}^{\mu})^{1/2}\,.$
Instead of relying on Lemma 3.3 (for which an analogue for $\operatorname{FI}_{\infty}$ is out of scope), suppose instead that for some constant $d_{T-t}\geq 0$
$\sqrt{\operatorname{FI}_{\infty}(q_{T-t}^{\nu}\,\|\,q_{T-t}^{\mu})}\leq d_{T-t}\sqrt{\operatorname{FI}_{\infty}(\nu\,\|\,\mu)}\,.$ (4.2)
Carrying out the rest of the argument as before, we obtain $\gamma$-a.s.
$\|X_{T}-Y_{T}\|\leq\sqrt{\operatorname{FI}_{\infty}(\nu\,\|\,\mu)}\,\eta_{T}\coloneqq\sqrt{\operatorname{FI}_{\infty}(\nu\,\|\,\mu)}\int_{0}^{T}d_{T-s}\exp\Bigl(\int_{0}^{T-s}\theta_{u}\,\mathrm{d}u\Bigr)\,\mathrm{d}s\,.$ (4.3)
It remains to show that (4.2) holds, after which taking limits in (4.3) would yield our desired result.
To this end, for $y\in\mathbb{R}^{d}$ and $t\geq 0$, we write
$\mu_{y,t}(x)\propto q_{t}(y\mid x)\,\mu(x)\,,\qquad q_{t}(\cdot\mid x)=\mathcal{N}(e^{-t}x,(1-e^{-2t})I)\,.$
Finally, we recall that $\mu$ satisfies a log-Sobolev inequality with constant $\lambda>0$ if
${\rm KL}(\nu\,\|\,\mu)\coloneqq\int\log\frac{\mu}{\nu}\,\mathrm{d}\mu\leq\frac{\lambda}{2}\,{\rm FI}(\nu\,\|\,\mu)\,.$
We will now prove the following.
Lemma 4.1.
Suppose that $\mu$ is such that for all $y\in\mathbb{R}^{d}$ and $t\geq 0$, $\mu_{y,t}$ satisfies a log-Sobolev inequality with constant $\lambda_{t}$. Then, for all $s\geq 0$,
$\sqrt{\operatorname{FI}_{\infty}(q^{\nu}_{s}\,\|\,q^{\mu}_{s})}\leq\frac{e^{s}\lambda_{s}}{e^{2s}-1}\sqrt{\operatorname{FI}_{\infty}(\nu\,\|\,\mu)}\,.$
Proof.  Recalling that $\rho Q_{t}(y)=\int q_{t}(y\mid x)\,\mathrm{d}\rho(x)$, it is easy to see that
$\nabla\log\frac{\nu Q_{s}}{\mu Q_{t}}(y)=\frac{e^{-s}}{1-e^{-2s}}\bigl(\mathbb{E}_{\nu_{y,s}}[X]-\mathbb{E}_{\mu_{y,s}}[X]\bigr)\,$,
and thus in norm
$\Bigl\|\nabla\log\frac{\nu Q_{s}}{\mu Q_{s}}\Bigr\|_{L^{\infty}(\nu Q_{s})}=\frac{e^{-s}}{1-e^{-2s}}\sup_{y\in\mathbb{R}^{d}}{\Bigl\lVert\int x\,\mathrm{d}(\nu_{y,s}-\mu_{y,s})(x)\Bigr\rVert}\,.$
We can further bound the right-hand side as
$\Bigl\|\int x\,\mathrm{d}(\nu_{y,s}-\mu_{y,s})(x)\Bigr\|\leq W_{2}(\nu_{y,s},\mu_{y,s})\,.$
If $\mu_{y,s}$ satisfies a log-Sobolev inequality with constant $\lambda_{s}$, it also satisfies the following transport-information inequality:
$W_{2}^{2}(\nu_{y,s},\mu_{y,s})\leq\lambda_{s}^{2}\,{\rm FI}(\nu_{y,s}\,\|\,\mu_{y,s})\,.$
(This is a generalization of (3.1); see [10].) By the definitions of $\mu_{y,s}$ and $\nu_{y,s}$,
$\nabla\log\frac{\nu_{y,s}}{\mu_{y,s}}=\nabla\log\frac{\nu}{\mu}\,.$
As $\sup_{y\in\mathbb{R}^{d}}{\rm FI}(\nu_{y,t}\,\|\,\mu_{y,t})\leq\operatorname{FI}_{\infty}(\nu\,\|\,\mu)$, our proof is complete. ∎
We now instantiate Lemma 4.1 for our existing examples to obtain contraction estimates in $\operatorname{FI}_{\infty}$.
Proposition 4.2.
For any $s\geq 0$, write $u(s)=e^{2s}-1$. If $\mu$ satisfies the conditions of
  1. (a)
    Corollary 3.5, then Lemma 4.1 holds with
    $\lambda_{s}=(\alpha+1/u(s))^{-1}\exp\Bigl(\frac{L^{2}}{\alpha+1/u(s)}+\frac{4L}{(\alpha+1/u(s))^{1/2}}\Bigr)\,;$ (4.4)
  2. (b)
    Corollary 3.7, then Lemma 4.1 holds with
    $\lambda_{s}=(1+\alpha+1/u(s))^{-1}\exp\Bigl(\frac{\hat{g}^{\prime}(0)}{1+\alpha+1/u(s)}\Bigr)\,.$ (4.5)
Proof.  Suppose $\mu$ satisfies the assumptions of Corollary 3.5, i.e., $\mu\propto\exp(-(V+H))$ where $V$ is $\alpha$-strongly convex and $H$ is $L$-Lipschitz. Then it is easy to see that for any $y\in\mathbb{R}^{d}$ and $s\geq 0$ that $\mu_{y,s}$ is strongly log-concave with parameter $\alpha+1/u(s)$ and the log-perturbation $H$ remains unchanged. Thus by [3, Theorem 1.4], $\mu_{y,s}$ satisfies the log-Sobolev inequality with parameter given by (4.4).
The argument for the second case is identical—the tilted measure is more strongly log-concave everywhere. In this case, the log-Sobolev constant is given by (4.5); see [8, Theorem 5.7].∎
Combined with the computations surrounding (4.2) and (4.3), we now arrive at the following result.
Theorem 4.3.
Suppose $\mu$ is such that ($\Theta$) holds for some constants $(\theta_{s})_{s\geq 0}$ and that $\mu_{y,t}$ satisfies a log-Sobolev inequality with constant $(\lambda_{t})_{t\geq 0}$ for any $y\in\mathbb{R}^{d}$ and $t\geq 0$. Then $\gamma$-a.s.,
$\|X_{T}-Y_{T}\|\leq\sqrt{\operatorname{FI}_{\infty}(\nu\,\|\,\mu)}\,\eta_{T}\coloneqq\sqrt{\operatorname{FI}_{\infty}(\nu\,\|\,\mu)}\int_{0}^{T}d_{T-s}\exp\Bigl(\int_{0}^{T-s}\theta_{u}\,\mathrm{d}u\Bigr)\,\mathrm{d}s\,$,
where $d_{T-s}=e^{T-s}\lambda_{u(T-s)}/u(T-s)$ and $u(T-s)\coloneqq e^{2(T-s)}-1$. Assuming that $\eta_{\infty}\coloneqq\lim_{T\to\infty}\eta_{T}$ exists, we have that
$\|T^{\mu}_{\rm KM}-T^{\nu}_{\rm KM}\|_{L^{\infty}(\gamma)}\leq\eta_{\infty}\sqrt{\operatorname{FI}_{\infty}(\nu\,\|\,\mu)}\,.$
We briefly mention that, to the best of our knowledge, Theorem 4.3 establishes a new transport-information inequality of the form
$W_{\infty}(\mu,\nu)\leq\|T^{\mu}_{\rm KM}-T_{\rm KM}^{\nu}\|_{L^{\infty}(\gamma)}\leq\eta_{\infty}\sqrt{\operatorname{FI}_{\infty}(\nu\,\|\,\mu)}\,$,
where we recall that $W_{\infty}(\mu,\nu)\coloneqq\inf_{\pi\in\Pi(\mu,\nu)}{\rm{ess\,sup}}_{(X,Y)\sim\pi}\|X-Y\|$, where $\Pi(\mu,\nu)$ is the set of joint measures with first- and second-marginal given by $\mu$ and $\nu$ respectively. In particular, we can revisit our previous examples (log-Lipschitz perturbations and strongly log-concave outside a ball) in the context of Theorem 4.3, where we obtain the same constants as Section 3.
Corollary 4.4.
Suppose $\mu\propto\exp(-(V+H))$ where $V$ is $\alpha$-strongly convex and $H$ is $L$-Lipschitz. Then Theorem 4.3 holds with constant
$\eta_{\infty}=\alpha^{-1}\exp\Bigl(\frac{3L^{2}}{2\alpha}+\frac{6L}{\sqrt{\alpha}}\Bigr)\,.$
Proof.  By Theorem 4.3, we can compute $d_{s}$ and use the bound $\alpha+1/u(s)\geq\alpha$ for all $s\geq 0$ to obtain
$d_{s}\leq\frac{e^{s}}{e^{2s}-1}\,\frac{1}{\alpha+(e^{2s}-1)^{-1}}\exp\Bigl(\frac{L^{2}}{\alpha}+\frac{4L}{\sqrt{\alpha}}\Bigr)=\frac{e^{s}}{\alpha\,(e^{2s}-1)+1}\exp\Bigl(\frac{L^{2}}{\alpha}+\frac{4L}{\sqrt{\alpha}}\Bigr)\,$,
which then leads to
$\eta_{T}\leq\exp\Bigl(\frac{L^{2}}{\alpha}+\frac{4L}{\sqrt{\alpha}}\Bigr)\int_{0}^{T}\frac{e^{(T-s)}}{\alpha\,(e^{2(T-s)}-1)+1}\exp\Bigl(\int_{0}^{T-s}\theta_{u}\,\mathrm{d}u\Bigr)\,\mathrm{d}s\,.$
We already computed the integral inside the exponential (in the proof of Corollary 3.5). Dropping the same terms, we obtain the following upper bound
$\eta_{T}\leq\exp\Bigl(\frac{3L^{2}}{2\alpha}+\frac{6L}{\sqrt{\alpha}}\Bigr)\int_{0}^{T}\frac{e^{2(T-s)}}{(\alpha e^{2(T-s)}+(1-\alpha))^{3/2}}\,\mathrm{d}s\,.$
One can obtain our final result by evaluating the integral by elementary means, and taking the limit as $T\to\infty$. ∎
Corollary 4.5.
Suppose $\mu\propto\gamma\exp(-h)$ where the potential satisfies the conditions in Corollary 3.7. Then Theorem 4.3 holds with constant
$\eta_{\infty}=(1+\alpha)^{-1}\exp\Bigl(\frac{3\hat{g}^{\prime}(0)}{2(1+\alpha)}\Bigr)\,.$
Proof.  The computations follow verbatim the arguments from Corollary 4.4, using the second part of Theorem 4.3, and are thus omitted. ∎
5 Conclusion
This note establishes, to our knowledge, the first stability properties of the Kim–Milman flow map. Some settings in our work (i.e., log-Lipschitz perturbations of strongly log-concave measures, or having asymptotically convex profiles) have not been analyzed for optimal transport maps or entropic transport maps. We show that our proof technique can be modified to give $L^{\infty}$-type stability bounds as well. In all cases, it is natural to extend these stability properties to measures which live on Riemannian manifolds, or to other flow maps (e.g., the flow map induced by general stochastic interpolants [1, 18] or the Schrödinger bridge flow map [15, 22, 25]).
Acknowledgments
We thank Zhou Fan and Daniel Lacker for helpful discussions. AAP thanks the Foundations of Data Science at Yale University for support. MSZ is funded by an NSERC CGS-D award.